An investment advisor invested $18000 in two accounts.One investment earned 11.2% annual simple interest while the other investment lost 4.7%. The total earnings from both investments were $1062. Find the amount invested at 11.2%.
I understand these interest problems, but I do not understand how to represent the loss of 4.7% in an equation. Your help would be greatly appreciated! Thank you.

Answer by mananth(16946)

Part I 11.20% per annum ------------- Amount invested =x
Part II -4.70% per annum ------------ Amount invested = y
18000
Interest----- 1062

Part I 11.20% per annum ---x
Part II -4.70% per annum ---y
Total investment
x + 1 y= 18000 -------------1
Interest on both investments
11.20% x + -4.70% y= 1062
Multiply by 100
11.2 x + -4.7 y= 106200.00 --------2
Multiply (1) by -11.2
we get
-11.2 x -11.2 y= -201600.00
Add this to (2)
0 x -15.9 y= -95400
divide by -15.9
y = 6000
Part I 11.20% $ 12000
Part II -4.70% $ 6000

CHECK
12000 --------- 11.20% ------- 1344.00
6000 ------------- -4.70% ------- -282.00
Total -------------------- 1062.00

m.ananth@hotmail.ca